# supplementary

### v1

@@ 5.3 Estimation error bound for linear regression
[…]
@@
[…]
@@ 5.4 Estimation error bound for logistic regression
Under Assumptions 5.2, 5.3, and 5.5, if we take
$\lambda=\frac{2\sqrt{2}C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{{\log(2p)}+{\log|\mathcal{I}|}},$
then with probability at least $1-2/p$,
$\|\hat{\beta}-\beta^{*}\|_{2}\;$ $\leq\;\frac{3\sqrt{2}C}{\kappa\sqrt{n}}\Theta(\mathcal{T})\sqrt{{\log(2p)}+{\log|\mathcal{I}|}}\sum_{\ell_{1}\in\mathcal{I}_{1}}\frac{1}{\sqrt{a_{\ell_{1}}}}+\frac{C}{\kappa}\sqrt{\frac{|\mathcal{N}|\log(2p|\mathcal{N}|)}{8n}}$ (5.4)
@@
6 Simulation
Various simulations are conducted to study the empirical properties of the proposed method. In the first subsection, experiments showcase the validity of our theory, emphasizing the impact of tree structure on prediction accuracy; The second subsection presents the comparison between our methods and several competitors, demonstrating a general advantage across different settings.
@@ 6.1 Validation of the theory
The theory we have built reveals two interesting facts: (i) The structure of the tree largely impacts the prediction and estimation accuracy, as observed in Theorem 5.1 and 5.2; (ii) Corollary 5.2 shows that, under some specific design of the tree, we are able to achieve a better prediction rate than that of the rare feature selection method (Yan and Bien, 2021). The two experiments in this subsection will show that these conclusions can also be drawn empirically.
@@ 6.1.1 Effect of number of internal nodes
We start from generating a fixed tree structure, as shown in Figure 3a. The total dimension $p$ of the predictors is set to be 48, represented by the leaves of the tree. We set the number of true effect groups $K=10$, the corresponding aggregating set is highlighted by those shaded latent nodes in the graph. The first 5 groups each has 9 features; and the remaining 5 groups each has 3 features. For such a given tree, We simulate $K=10$ independent values from Uniform(1.5,2.5) and alternate the signs to obtain the true effect values. The entries of the design matrix $X\in\mathbb{R}^{n\times 60}$ are generated independently from Poisson(0.02). Then we obtain $y\sim N(X\beta^{*},\sigma^{2}\mathbf{I}_{n})$, where $\sigma^{2}=\|X\beta^{*}\|_{2}^{2}/(5n)$. We train the model on a training set with size 50 and select the penalty parameter on a validation set with the same sample size. The prediction errors are evaluated on an independent validation set with size 500. We repeat the procedure 200 times and apply the proposed method to the generated datasets.
@@
Figure 3: Results of simulation 6.1.1. (a) The original tree structure $\mathcal{T}_{0}$, from which we generate $\mathcal{T}_{1},\cdots,\mathcal{T}_{5}$ by deleting nodes; (b) The boxplot of the prediction error of the four methods under 200 replications; (c) The contrast of group accuracy of our method and RARE under different tree structures.
@@
The first column in Figure 3a shows our baseline behavior, using the full tree and compared to the rare feature selection method. Now we modify the tree structure by randomly deleting latent nodes, either above or below the aggregating set. However, we preserve the entire aggregating set so that Assumption 5.2 still holds. Upon deletion of a node, we connect all of its children directly to its parent. We use $\mathcal{T}_{0}$ to denote the full tree, and generate trees $\mathcal{T}_{1},\cdots,\mathcal{T}_{5}$ by deleting 10, 20, 30 nodes below and 3, 6 nodes above the aggregating set, respectively. Figure 3b and 3c summarize the predicted error and adjusted Rand index using the same datasets but different trees. A dramatic increase in the prediction error and decrease in the grouping accuracy are observed when more latent nodes are deleted below the aggregating set, where the corresponding parameters are supposed to be merged. In contrast, deleting nodes above the aggregating set improves the behavior and gives a smaller prediction error and a larger Rand index. It aligns with the theorem that we prefer a tree with as much precise information as possible above the true aggregating set; and as many branches as possible below the aggregating set, to reduce the volume of the restricted set $S_{\mathcal{T}}$ and obtain a larger strong convexity constant $\kappa$. We also measure the prediction error using oracle least square and oracle ridge as the benchmark performance. They assume we have the oracle group information and directly regress over the correctly aggregated groups. With a proper tree structure such as $\mathcal{T}_{5}$, we come close to the most ideal situation.
@@ 6.1.2 Effect of tree size
Next, we simulate the scenario in Example 2, Section 5.2. The number of total groups $K$ is assumed to be fixed, and we evaluate the performance of the methods when the dimension $p$ increases. Figure 4a shows the metastructure of our generated trees, with the leaves representing the aggregating set of size $K=20$. Each node in the aggregating set is the root of a replication for a subtree with $p_{s}$ dimension of features. Therefore, the total dimension $p=20p_{s}$ and the growing structure below the aggregating set are ignored in the plot. In the simulation, we choose $p_{s}\in\{3,5,10,20,30,40,50\}$, respectively, so that $p\in\{60,100,200,400,500,800,1000\}$. We train the model on a training set with size 50, and choose the penalty parameter using the square loss on a validation set with the same sample size. The prediction error for our method and the rare feature selection method is evaluated on an independent test set with size 500.
Figure 4: Results of experiment 6.1.2. (a) The meta-structure from which we generate trees by adding nodes below the colored ones; (b) The boxplot of the ratio between our and RARE prediction errors under 200 replications, the curve fits the average ratio versus $1/\sqrt{\log(p)}$; (c) The contrast of group accuracy of our method and RARE under different tree structures and dimensions.
On the theory side, Corollary 5.2 told us that we should have a finite sample rate $1/\sqrt{\log(p)}$ times faster than the rare feature selection method. Though it is merely a description of the error bound, which we are actually unsure about how tight it could be, we can still observe the effect empirically. Figure 4b depicts the ratio between prediction errors of our method and the rare method, in 200 replications. We denote the mean value of each setting using the white dots and fit a linear function versus $1/\sqrt{\log(p)}$. The white line plots the predicted curve $r=1.0659/\sqrt{\log(p)}+0.2145$, with the coefficient significant under the level 0.05. The decreasing trend in the ratio is evident and aligns with the theory. It can also be observed that the ratio is mostly below 1, indicating a lower prediction error than the competitor. In addition, we have the extra bonus of better group selection accuracy, as shown in Figure 4c.
@@ 6.2 Performance in general settings
[…]
Figure 5: Results of the three Scenarios. The first row gives the prediction error, and the second row gives the group selection accuracy.
Our advantage in group selection consistency, although not shown theoretically, can be heuristically explained. The rare feature selection method, as a LASSO-type method, does variable selection on the synthetic predictors created by summations over the real predictors under each latent node. The status of each node is decided ”independently” from the corresponding synthetic predictor. Therefore, for a specific group to be precisely aggregated, all the ancestors of the group latent node have to be simultaneously shrunk to 0, which is generally hard considering the number of nodes. In contrast, our method only requires that the group standard deviation of the corresponding latent node be shrunk to 0. The differences, reflected by a much simpler and smoother solution path, are visualized in the following figures. Figure 6 plots the solution paths of our method and RARE, under the full tree $\mathcal{T}_{0}$ in Experiment 1.
Figure 6: Solution Paths Comparison
In addition, LASSO is well known to be largely affected by the magnitude of predictors in the design matrix, but RARE is not able to standardize before the variable selection step (otherwise, the aggregation is improper). Thus, it tends to over-penalize the predictors with smaller magnitude. There arises an unfair selection problem between synthetic predictors that are made up of more siblings and fewer siblings within the tree, and it gets more severe when the tree is highly imbalanced. This property also contributes to the less satisfactory grouping accuracy of RARE and is further illustrated in the real data analysis.
@@ 7 Microbiome data analysis
We apply the proposed method to analyze microbiome data. As discussed in introduction, we aim to simultaneously choose the aggregation level of OTUs and estimate their group effects specific to each biomedical trait. In this subsection, we provide a procedure for learning and testing group effects using our proposed method, with examples demonstrating our advantage in scientific discovery.
7.1 Prediction
We consider the dataset from Sinha et al. (2016), where researchers collected fecal samples from a total of 131 individuals with colorectal cancer (CRC) and healthy controls. They profiled the microbial communities of the samples (through 16S rRNA gene sequencing) up to the genus level and summarized with a compositional matrix $X\in\mathbb{R}^{131\times 85}$. A taxonomic tree, as shown in Figure 7, describes the biological relations between the 85 genes. We attempt to predict the continuous outcome, the body mass index (BMI) of each individual; and the binary outcome, the positivity of CRC, respectively. We adopt the relative shift model proposed by Li et al. (2023), which works directly with the compositional predictors and omits the intercept.
$E[y_{i}|X_{i}]=f^{-1}(\beta_{1}X_{i1}+\cdots+\beta_{p}X_{ip}).$ (7.1)
Compared to the conventional log-contrast model (Aitchison and Bacon-Shone, 1984), the relative-shift model directly accommodates zero-valued predictors, and enjoys the property of scale and shift invariance. Each contrast between distinct coefficients $\beta_{i}-\beta_{j}$ is interpreted as the relative shift effect between two groups. The model then allows meaningful aggregation of the predictors (Li et al., 2023), it therefore fits perfectly with our framework. We use the squared loss for modeling the BMI, and cross-entropy loss for the CRC outcome.
@@ 7.1 Prediction
Sinha et al. (2016) BMI 131×85131$\times$ 85131 × 85 0.281 18.216 18.669 37.813 118.679
Erawijantari et al. (2020) BMI 96×192196$\times$ 192196 × 1921 0.627 9.130 9.312 549.728 160.635
Wang et al. (2020) BMI 287×2853287$\times$ 2853287 × 2853 0.415 16.034 16.027 46.598 531.252
Yachida et al. (2019) BMI 347×3418347$\times$ 3418347 × 3418 0.274 10.509 10.391 2328.520 209.242
@@
Sinha et al. (2016) CRC 131×85131$\times$ 85131 × 85 0.566/0.576 0.637 0.632 0.625 0.635
Kim et al. (2020) CRC 240×121240$\times$ 121240 × 121 0.514/0.522 0.690 0.692 0.698 0.690
Jacobs et al. (2016) IBD 90×50790$\times$ 50790 × 507 0.653/0.582 0.667 0.704 0.622 0.755
Erawijantari et al. (2020) GC* 96×192196$\times$ 192196 × 1921 0.551/0.539 0.704 0.698 0.726 0.730
Franzosa et al. (2019) IBD* 220×2834220$\times$ 2834220 × 2834 0.801/0.822 0.472 0.453 0.474 0.553
Wang et al. (2020) ESRD* 287×2853287$\times$ 2853287 × 2853 0.834/0.832 0.423 0.419 0.376 0.687
Yachida et al. (2019) CRC 347×3418347$\times$ 3418347 × 3418 0.548/0.450 0.662 0.661 0.686 0.671
Table 1: Results on other microbiome datasets.
The patterns we discover here turn out to hold true for many other datasets with even higher dimensions, as summarized in the following Table 1. The table lists the sources of the datasets, along with their outcome types, dimensions, estimated signal strengths, and the prediction error means among 200 resamples. The patterns align with the findings of Bien et al. (2021), where there is generally not much improvement in prediction using the RARE for microbiome datasets. As was mentioned in both Bien et al. (2021) and Bichat et al. (2020), whether it is useful to incorporate taxonomic information into statistical modeling is still an ongoing debate. Our method and the analysis for signal strength also serve to provide a new perspective on this problem.
Figure 7: Aggregated results
@@ 7.2 Post-selection inference on aggregated effects
To avoid double dipping and conduct post-selection inference, people often need to split the original dataset $S=\{(X_{1},y_{1}),\cdots,(x_{n},y_{n})\}$ into two parts for model selection and inference, respectively. With the binary outcome CRC, we apply the data fission method proposed by Neufeld et al. (2024). Instead of splitting data points into two independent parts, Leiner et al. (2023) proposed data fission that aims to partition the information and create two datasets each of the original sample size. Specifically, the set $S$ undergoes fission to form two datasets $S_{1}=\{(X_{1},y_{1}^{(1)}),\cdots,(x_{n},y_{n}^{(1)})\}$ and $S_{2}=\{(X_{1},y_{1}^{(2)}),\cdots,(x_{n},y_{n}^{(2)})\}$, both with the same design matrix, but have outcomes randomly perturbed in different ways. After fission, the two data sets should together be able to exactly recover the original data set so that no information is lost. Neufeld et al. (2024) improves the methodology under logistic regression setting and allows inference on the exact parameter target. Under the assumed model
$y_{i}$ $\sim\text{Bernoulli}(\theta_{i}),$ (7.2)
$\theta_{i}$ $=\frac{\exp(X_{i}^{T}\beta)}{1+\exp(X_{i}^{T}\beta)},$ (7.3)
we generate the perturbations $Z_{i}\sim\text{Bernoulli}(\delta)$ i.i.d. for $i=1,\cdots,n$, such that the partitioned points $y^{(1)}_{i}=(1-Z_{i})y_{i}+Z_{i}(1-y_{i})$ and $y^{(2)}_{i}=y_{i}.$ Their new distributions and log-odds ratio are
$y^{(i)}_{1}$ $\sim\text{Bernoulli}(\theta_{i}+\delta-2\theta_{i}\delta),$ (7.4)
$y^{(2)}_{i}|y^{(1)}_{i}$ $\sim\text{Bernoulli}\left(\frac{\theta_{i}}{\theta_{i}+(1-\theta_{i})(\frac{\delta}{1-\delta})^{2y^{(1)}_{i}-1}}\right),$ (7.5)
$\log\left(\frac{\mathop{\rm Pr(y_{i}^{(2)}=1|y_{i}^{(1)})}}{\mathop{\rm Pr(y_{i}^{(2)}=0|y_{i}^{(1)})}}\right)$ =$=\begin{cases}\log\left(\frac{1-\delta}{\delta}\right)+X_{i}^{T}\beta,&\text{if }y^{(1)}_{i}=1,\\ \log\left(\frac{\delta}{1-\delta}\right)+X_{i}^{T}\beta,&\text{if }y^{(1)}_{i}=0.\end{cases}$ (7.6)
It now allows one to perform model selection on the first set $S_{1}$, and conduct inference on the second set $S_{2}$ conditioned on all information in $S_{1}$. To infer on the parameter $\beta$, we only need to fit the standard logistic GLM on $S_{2}$ with the offset depending on $\delta$ as shown in (7.6). The parameter $\delta$ controls the proportion of information assigned to $S_{1}$ and $S_{2}$. It varies from 0.5 to 1 when increasing information is assigned to the selection set $S_{1}$, data fission thus can be viewed as a continuous counterpart to data splitting. Meanwhile, it inherits the benefit of data carving (Fithian et al., 2014) that allows all data points to participate in both selection and inference, thus always conditions on the same design matrix $X$ and mitigates the potential problem caused by the imbalanced influences of different data points.
@@
In our case, we apply the data fission method to the CRC outcome with $\delta=0.9$, placing a larger fraction of the information in the selection set. After obtaining the fissured sets, we perform model selection on $S_{1}$ using a 5-fold cross-validation, for LASSO, RARE and our method, respectively. Figure 7 demonstrates the aggregated predictor groups using different colors. RARE aggregates the predictors into two groups, much fewer than the twelve groups detected by our method. Although the quantities of groups might not have a privilege over one another, the problem is that RARE, not being a group-wise penalty, violates the tree structure to obtain the highly aggregated groups.
@@
We then select or aggregate the predictors in $S_{2}$ based on the models selected in the first step, and fit the standard logistic regression using the offsets provided as in (7.6). Inference is carried out on the fitted model for those aggregated or selected predictors, whose fitted values and significance are also labeled in Figure 7. While both RARE and our method detect one significant taxonomic group Clostridia, we have another significant feature, Fusobacterium at Genus level, that is overlooked by RARE. The OTU Fusobacterium is not only identified by the LASSO procedure to be statistically significant, but also a well-studied type of bacteria that is highly associated with CRC in biomedical studies(Kostic et al., 2013). We also conduct a multiple testing on the effect contrasts between Fusobacterium and the other eleven aggregated gorups. The p-values, after Benjamini-Hochberg correction procedure(Benjamini and Hochberg, 1995), are shown in Table 2 to be all significant. The OTU Fusobacterium is noticably rare within the sample, only 17 out of 131 individuals have a nonzero value on this feature. Meanwhile, the tree is highly imbalanced, not allowing it to be aggregated with any other predictors. As a result, it participates in variable selection with an overly small magnitude in the design matrix, and is therefore screened out by RARE to be aggregated with another group.

### v2

@@ 5.3 Estimation error bound for linear regression
$S_{\mathcal{T}}\;=\;\Bigl\{\Delta:\;\sum_{\ell_{0}\in\mathcal{I}_{0}}\frac{1}{\sqrt{a_{\ell_{0}}}}\,\|M({\mathcal{A}_{\ell_{0}}})\Delta\|_{2}\;\leq\;3\sum_{\ell_{1}\in\mathcal{I}_{1}}\frac{1}{\sqrt{a_{\ell_{1}}}}\,\|M({\mathcal{A}_{\ell_{1}}})\Delta\|_{2}\;+\;\frac{\sqrt{|\mathcal{N}|}}{8\Theta(\mathcal{T})}\|\Delta\|_{2}\Bigr\}.$
@@
Under Assumptions 5.1, 5.2, 5.3, and the restricted eigenvalue condition 5.4, if the tuning parameter $\lambda$ is chosen as in Theorem 5.1, then with probability at least 12/p1-2/p,
@@ 5.4 Estimation error bound for logistic regression
Under Assumptions 5.2, 5.3, and 5.5, if we take $\lambda=\frac{2\sqrt{2}C}{\sqrt{n}}\Theta(\mathcal{T})\sqrt{{\log(2p)}+{\log|\mathcal{I}|}},$ then with probability at least 12/p1-2/p,
$\|\hat{\beta}-\beta^{*}\|_{2}\;$ $\leq\;\frac{3\sqrt{2}C}{\kappa\sqrt{n}}\Theta(\mathcal{T})\sqrt{{\log(2p)}+{\log|\mathcal{I}|}}\sum_{\ell_{1}\in\mathcal{I}_{1}}\frac{1}{\sqrt{a_{\ell_{1}}}}+\frac{C}{\kappa}\sqrt{\frac{|\mathcal{N}|\log(2p|\mathcal{N}|)}{8n}}\vskip-10.0pt$ (5.4)
@@
6 Simulation studies
@@ 6.1 Validation of the theory
The theory we established reveals two interesting facts: (i) the structure of the tree largely impacts the prediction and estimation accuracy, as observed in Theorem 5.1 and 5.2; (ii) under some specific design of the tree, we are able to achieve a better prediction rate than that of the method of Yan and Bien (2021) (Corollary 5.2). The two experiments in this subsection show that these conclusions can also be drawn empirically.
@@ 6.1.1 Effect of number of internal nodes
We start from generating a fixed tree structure, as shown in Figure 3a. The total dimension $p$ of the predictors is set to be 60, represented by the leaves of the tree. We set the number of true effect groups $K=10$, the corresponding aggregating set is highlighted by those shaded latent nodes in the graph. The first 5 groups each has 9 features; and the remaining 5 groups each has 3 features. For such a given tree, We simulate $K=10$ independent values from Uniform(1.5,2.5) and alternate the signs to obtain the true effect values. The entries of the design matrix $X\in\mathbb{R}^{n\times 60}$ are generated independently from Poisson(0.02). Then we obtain $y\sim N(X\beta^{*},\sigma^{2}\mathbf{I}_{n})$, where $\sigma^{2}=\|X\beta^{*}\|_{2}^{2}/(5n)$. We train the model on a training set with size 50 and select the penalty parameter on a validation set with the same sample size. The prediction errors are evaluated on an independent validation set with size 500. We repeat the procedure 200 times and apply the proposed method to the generated datasets.
@@
Figure 3: Results of the simulation studies described in Section 6.1.1. (a) The original tree structure $\mathcal{T}_{0}$, from which we generate $\mathcal{T}_{1},\cdots,\mathcal{T}_{5}$ by deleting nodes. (b) The boxplot of the prediction error of the four methods under 200 replications. (c) The contrast of group accuracy (adjusted Rand index) of our method and RARE under different tree structures.
@@
The first column in Figure 3b shows our baseline behavior, using the full tree and compared to the rare feature selection method. Now we modify the tree structure by randomly deleting latent nodes, either above or below the aggregating set. However, we preserve the entire aggregating set so that Assumption 5.2 still holds. Upon deletion of a node, we connect all of its children directly to its parent. We use $\mathcal{T}_{0}$ to denote the full tree, and generate trees $\mathcal{T}_{1},\cdots,\mathcal{T}_{5}$ by randomly deleting 10, 20, 30 nodes below (in $\mathcal{I}_{0}$) and 3, 6 nodes above the aggregating set (in $\mathcal{I}_{1}$), respectively. See the Supplementary Materials LABEL:appendix.tree.plots for a visualization of the tree structures. Figure 3b and 3c summarize the predicted error and adjusted Rand index using the same datasets but different trees. A dramatic increase in the prediction error and decrease in the grouping accuracy are observed when more latent nodes are deleted below the aggregating set, where the corresponding parameters are supposed to be merged. In contrast, deleting nodes above the aggregating set improves the behavior and gives a smaller prediction error and a larger Rand index. It aligns with the theorem that we prefer a tree with as much precise information as possible above the true aggregating set; and as many branches as possible below the aggregating set, to reduce the volume of the restricted set $S_{\mathcal{T}}$ and obtain a larger strong convexity constant $\kappa$. We also measure the prediction error using oracle least square and oracle ridge as the benchmark performance. They assume we have the oracle group information and directly regress over the correctly aggregated groups. With a proper tree structure such as $\mathcal{T}_{5}$, we come close to the most ideal situation.
@@ 6.1.2 Effect of tree size
Next, we simulate data under the model described in Example 2, Section 5.2. The number of total groups $K$ is assumed to be fixed, and we evaluate the performance of the methods when the dimension $p$ increases. Figure 4a shows the metastructure of our generated trees, with the leaves representing the aggregating set of size $K=20$. Each node in the aggregating set is the root of a replication for a subtree with $p_{s}$ dimension of features. Therefore, the total dimension $p=20p_{s}$ and the growing structure below the aggregating set are ignored in the plot. In the simulation, we choose $p_{s}\in\{3,5,10,20,30,40,50\}$, respectively, so that $p\in\{60,100,200,400,500,800,1000\}$. We train the model on a training set with size 50, and choose the penalty parameter using the square loss on a validation set with the same sample size. The prediction error for our method and the rare feature selection method is evaluated on an independent test set with size 500.
Figure 4: Results of simulation studies described in Section 6.1.2. (a) The meta-structure from which we generate trees by adding nodes below the colored ones. (b) The boxplot of the ratio between our and RARE prediction errors under 200 replications, the curve fits the average ratio versus $1/\sqrt{\log(p)}$. (c) The contrast of group accuracy of our method and RARE under different tree structures and dimensions.
On the theory side, Corollary 5.2 suggests that we should have a finite sample rate $\{\log(p)\}^{-1/2}$ times faster than the rare feature selection method. Though it is merely a description of the error bound, which we are actually unsure about how tight it could be, we can still observe the effect empirically. Figure 4b depicts the ratio between prediction errors of our method and the rare method, in 200 replications. We denote the mean value of each setting using the white dots and fit a linear function versus $\{\log(p)\}^{-1/2}$. The black line plots the predicted curve $r=1.0659/\sqrt{\log(p)}+0.2145$, with the coefficient significant under the level 0.05. The decreasing trend in the ratio is evident and aligns with the theory. It can also be observed that the ratio is mostly below 1, indicating a lower prediction error than the competitor. In addition, we have the extra bonus of better group selection accuracy, as shown in Figure 4c.
@@ 6.2 Performance in general settings
  • Scenario 3: The outcome $y$ is binary. We set $n=50$, $K=20$, and allow dimension to grow with $p=\{50,100,200,400,600,800,1000\}$. $y$ is generated independently from $y_{i}\simBernoulli$$(1/(1+\exp\{-X_{i}^{T}\beta^{*}\}))$.
In each scenario, we select the penalty parameter on a validation set with the same sample size $n$ as the training set. The prediction error is then evaluated on an independent test set with size $10n$. The procedure is repeated 200 times, and Figure 5 summarizes the performances. The results indicate a general advantage in prediction, especially under high-dimensional settings. Scenario 3 further validates the proposed method for classification.
Figure 5: Simulation study results under the three scenarios described in Section 6.2. The first row gives the prediction error, and the second row gives the group selection accuracy measured by adjusted Rand index.
Our advantage in group selection consistency, although not shown theoretically, can be heuristically explained. The rare feature selection method, as a LASSO-type method, does variable selection on the synthetic predictors created by summations over the real predictors under each latent node. The status of each node is decided individually from the corresponding synthetic predictor. Therefore, for a specific group to be precisely aggregated, all the ancestors of the group latent node have to be simultaneously shrunk to zero, which is generally hard considering the number of nodes. In contrast, our method only requires that the group standard deviation of the corresponding latent node be shrunk to zero. The differences are reflected by a much simpler and smoother solution path (Section LABEL:appendix.solution.path, Supplementary Materials).
In addition, LASSO is well known to be largely affected by the magnitude of predictors in the design matrix, but RARE is not able to standardize before the variable selection step (otherwise, the aggregation is improper). Thus, RARE tends to over-penalize the predictors with smaller magnitude. Therein arises an unfair selection problem between synthetic predictors that are made up of more siblings and fewer siblings within the tree, and it gets more severe when the tree is highly imbalanced. This property also contributes to the less satisfactory grouping accuracy of RARE and is illustrated in the real data analyses.
@@ 7 Microbiome data analysis
We apply the proposed method to analyze microbiome data. As discussed in introduction, we aim to simultaneously choose the aggregation level of OTUs and estimate their group effects specific to each biomedical trait. In this section, we provide a procedure for learning and testing group effects, with examples demonstrating our advantage in scientific discovery.
7.1 Prediction accuracy under relative-shift model
We consider the dataset from Sinha et al. (2016), where researchers collected fecal samples from a total of 131 individuals with colorectal cancer (CRC) and healthy controls. They profiled the microbial communities of the samples (through 16S rRNA gene sequencing) up to the genus level and summarized with a compositional matrix $X\in\mathbb{R}^{131\times 85}$. A taxonomic tree, as shown in Figure 6, describes the biological relations between the 85 genes. We attempt to predict the continuous outcome, the body mass index (BMI) of each individual; and the binary outcome, the positivity of CRC, respectively. We adopt the relative shift model proposed by Li et al. (2023), which works directly with the compositional predictors and omits the intercept, i.e., the model assuming ${\rm E}[y_{i}|X_{i}]=f^{-1}(\beta_{1}X_{i1}+\cdots+\beta_{p}X_{ip}).$ Compared to the conventional log-contrast model (Aitchison and Bacon-Shone, 1984), the relative-shift model directly accommodates zero-valued predictors, and enjoys the property of scale and shift invariance. Each contrast between distinct coefficients $\beta_{i}-\beta_{j}$ is interpreted as the relative shift effect between two groups. The model then allows meaningful aggregation of the predictors (Li et al., 2023), it therefore fits perfectly with our framework. We use the squared loss for modeling the BMI, and cross-entropy loss for the CRC outcome.
@@ 7.1 Prediction accuracy under relative-shift model
Sinha et al. (2016) BMI $131\times 85$ 0.281 18.216 18.669 37.813 118.679
Erawijantari et al. (2020) BMI $96\times 1921$ 0.627 9.130 9.312 549.728 160.635
Wang et al. (2020) BMI $287\times 2853$ 0.415 16.034 16.027 46.598 531.252
Yachida et al. (2019) BMI $347\times 3418$ 0.274 10.509 10.391 2328.520 209.242
@@
Sinha et al. (2016) CRC $131\times 85$ 0.566/0.576 0.637 0.632 0.625 0.635
Kim et al. (2020) CRC $240\times 121$ 0.514/0.522 0.690 0.692 0.698 0.690
Jacobs et al. (2016) IBD $90\times 507$ 0.653/0.582 0.667 0.704 0.622 0.755
Erawijantari et al. (2020) GC* $96\times 1921$ 0.551/0.539 0.704 0.698 0.726 0.730
Franzosa et al. (2019) IBD* $220\times 2834$ 0.801/0.822 0.472 0.453 0.474 0.553
Wang et al. (2020) ESRD* $287\times 2853$ 0.834/0.832 0.423 0.419 0.376 0.687
Yachida et al. (2019) CRC $347\times 3418$ 0.548/0.450 0.662 0.661 0.686 0.671
Table 1: Prediction results on multiple microbiome datasets with either continuous or binary outcomes.
The patterns we discover here turn out to hold true for many other datasets with even higher dimensions, as summarized in Table 1. The table lists the sources of the datasets, along with their outcome types, dimensions, estimated signal-to-noise ratios, and the prediction error means among 200 resamples. The patterns align with the findings of Bien et al. (2021), where there is generally not much improvement in prediction using the RARE for microbiome datasets. As was mentioned in both Bien et al. (2021) and Bichat et al. (2020), whether it is useful to incorporate taxonomic information into statistical modeling is still an ongoing debate. In addition, a discussion of the impact of signal-to-noise ratio is provided in Supplementary Material LABEL:appendix.signal-noise.
@@ 7.2 Post-selection inference on aggregated effects
To avoid double-dipping and conduct post-selection inference, it is common to split the original dataset $S=\{(X_{1},y_{1}),\cdots,(X_{n},y_{n})\}$ into two parts for model selection and inference, respectively. With the binary outcome CRC, we apply the data fission method proposed by Neufeld et al. (2024). Instead of splitting data points into two independent parts, Leiner et al. (2023) proposed data fission that aims to partition the information and create two datasets each of the original sample size. Specifically, the set $S$ undergoes fission to form two datasets $S_{1}=\{(X_{1},y_{1}^{(1)}),\cdots,(X_{n},y_{n}^{(1)})\}$ and $S_{2}=\{(X_{1},y_{1}^{(2)}),\cdots,(X_{n},y_{n}^{(2)})\}$, both with the same design matrix, but have outcomes randomly perturbed in different ways. After fission, the two data sets should together be able to exactly recover the original data set so that no information is lost. In the sparse design context, data fission enjoys a unique advantage of stability by preserving the entire design matrix in both the training and inference sets. Neufeld et al. (2024) improves the methodology under the logistic regression setting and allows inference on the exact parameter target, with detailed implementation in Supplementary Material Section LABEL:appendix.fission.
@@
After obtaining the fissured sets, we perform model selection on $S_{1}$ using 5-fold cross-validation for LASSO, RARE, and our method. Figure 6 demonstrates the aggregated predictor groups using different colors. RARE aggregates the predictors into two groups, much fewer than the twelve groups detected by our method. Although the quantities of groups might not have a privilege over one another, the problem is that RARE, not being a group-wise penalty, violates the tree structure to obtain the highly aggregated groups.
We then select or aggregate the predictors in $S_{2}$ based on the models selected in the first step, and fit the standard logistic regression using the offsets provided as in (LABEL:mid.log-odds). Inference is carried out on the fitted model for those aggregated or selected predictors, whose fitted values and significance are also labeled in Figure 6.
Figure 6: Visualization of the fitted models and results of the post-selection inference procedure based on (a) using our proposed method and (b) RARE (Yan and Bien, 2021).
@@
While both RARE and our method detect one significant taxonomic group Clostridia, we have another significant feature, Fusobacterium at Genus level, that is overlooked by RARE. The OTU Fusobacterium is not only identified by the LASSO procedure to be statistically significant, but also a well-studied type of bacteria that is highly associated with CRC in biomedical studies (Kostic et al., 2013). We also conduct a multiple testing on the effect contrasts between Fusobacterium and the other eleven aggregated gorups. The p-values, after Benjamini-Hochberg correction procedure(Benjamini and Hochberg, 1995), are shown in Table 2 to be all significant. The OTU Fusobacterium is noticably rare within the sample, only 17 out of 131 individuals have a nonzero value on this feature. Meanwhile, the tree is highly imbalanced, not allowing it to be aggregated with any other predictors. As a result, it participates in variable selection with an overly small magnitude in the design matrix, and is therefore screened out by RARE to be aggregated with another group.
@@
Data Availability Statement
The authors confirm that the data supporting the findings of this study are available in its supplementary materials, and at https://github.com/JinwenFu001/direct_feature_aggregation.